OpenStudy (anonymous):

Let u = <-5, 1>, v = <7, -4>. Find 9u - 6v.

OpenStudy (jhannybean):

What is \(\langle -5 \cdot 9 ~,~ 1 \cdot 9\rangle -\langle 7 \cdot 6 ~,~ -4 \cdot 6\rangle\)?

OpenStudy (anonymous):

<-45, 9> - <42, -24>

OpenStudy (anonymous):

is it <-87, 33> ?

OpenStudy (jhannybean):

Okay, now we just subtract components from both u and v appropriately. \[\langle -45 -42 ~,~ 9 - (-24)\rangle\]

OpenStudy (jhannybean):

You got it

OpenStudy (anonymous):

thank you!!

OpenStudy (jhannybean):

Np :)